OpenStudy (anonymous):

g(x) = 2(x+1)^3 +8 (condition: x>1) obtain an expression for g'(x) and use your answer to explain why g(x) has an inverse.

OpenStudy (anonymous):

chain rule g'(x)=h(j(x))'j'(x) h(x)=2x^3, j(x)=x+1, h'(x)=6x^2, j'(x)=1 6*(x+1)^2 6(x^2+2x+1) If x>1 then g'(x) is always >0, then it's strictly increasing. As g(x) is strictly increasing to all x>1, then it's inyective. As g(x) is never 0, then g^-1(x) is defined to all x>1